solve the equation
(D2-3D+2)y=sin(ex) wherw D =D/DX

Grade:12th pass

1 Answers

Abhishek Sharma
4 years ago
By finding c.f and p.i we can get the solution of this differential equation.
Cf=c1(e^ax)+c2(e^bx) where a,b are roots of D^2-3D+2
IN p.i put D^2=-a^2